OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8 A. one solution B. two solutions C. infinitely many solutions D. no solution

OpenStudy (anonymous):

C when you simplify the first equation you get x=-4(y+1) and when you simplify the second one by dividing by 2 you get x+4y=4 or x=-4y+4 which gives x=-4(y+1) as well